40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 3 of 59  Not logged in ELOG logo
ID Date Author Type Category Subject
  2836   Fri Dec 17 12:46:41 2021 aaronSummaryPSOMAPSOMA calibration parameters, minor modifications

I removed the 10 dB attenuator between the LB box output and the laser driver's current modulation input. The ranges of the LB box output and ITC 502 input are matched, so there should be no need for an attenuator. I also slightly reduced the servo gain to 5.3. There is still a 10dB attenuator between the PDH mixer/lowpass and the LB box input, to avoid saturating the servo's input stage (+- 330 mV).

After moving most electronics to the same power strip as in the above elog and tying the table to ground, I no longer see a 144 kHz oscillation in the PDH error signal (oscilloscope). I do see a ~40 Hz oscillation in the PDH error signal... could that be the fundamental mode of the cantilever? Zach's thesis quotes a fundamental resonance at 45 Hz. The 40 Hz oscillation is present even while unlocked (but with the servo engaged).

Though the error signal looked clean, the lock was not steady for more than 10s of seconds. I suspected that the current control loop was far from the center of its range. I am monitoring the current control signal from the LB box Aux output, which picks off the signal before the final summing amplifier stage. Therefore, in order to adjust the temperature setpoint to 'zero' the control signal, I needed to first disconnect the LB box inputs then adjust the output offset until the voltage at aux output matched that at the final output. Then, with the laser locked, I adjusted the temperature setpoint until the control signal as measured at aux output was zero (plus some offset that was present at both the main output and aux output even with the inputs disconnected).

After adjusting the temperature setpoint manually, the beamspot is noticeably brighter on the trans mon camera, and the lock is stable for minutes.

I tried setting up the digital control to handle the temperature loop, but the realtime system is in a bad state and I didn't attempt to diagnose it.

  2835   Thu Dec 16 17:07:24 2021 aaronDailyProgressLaserbeat note setup

confirming fiber path

I traced the fiber beam path and confirmed that the system is set up as in elog 2833. I add a FC/PC-to-FC/APC patch cable between the 50-50 beamsplitter and the 1611FC-AC photodiode. We've been meaning to add that patch for a while to avoid damaging the connection.

I'd like to also see a beat note so we can hunt for this 140 kHz oscillation tomorrow.

I opened the connector housings on the cable for the N laser TEC and readjusted them such that the internal clamp bites down on the jacket of the cable, rather than closing over the conductor wires themselves (see attachments 1 and 2). Wiggling the cable no longer causes an alarm on the TEC.

I noticed that the FC 1611 photodiode was being powered from a GlobTek +-12 V DC supply with 0.25A current limit; however, those photodiodes require +-15 V DC with 0.3A current limit. I located the proprietary cable and am now powering that photodiode from the New Focus 0901 power supply located on the PSOMA rack.

I also noticed that the SMA cable from the fiber box panel to the spectrum analyzer that should have been carrying the beat note was connected to the wrong port! That's probably why it wasn't visible before.

With those changes, I could see the beat note on the spectrum analyzer yes.

I wanted to add a cable from the beat note PD DC output to the fiber box panel, so I could check that the DC power on that photodiode makes sense. But, I couldn't locate one in cryo or EE. Maybe the 40m or another WB lab has a spare?

Attachment 1: loose.jpg
loose.jpg
Attachment 2: clamped.jpg
clamped.jpg
  2834   Tue Dec 14 14:10:36 2021 aaronSummaryPSOMAPSOMA calibration parameters, minor modifications

labwork to do

  • posting table of important parameters (like those in elog 2830) in wiki, to be printed and posted in the lab
  • psoma servo drawing

done

  • AC power for (most) electronics involved in locking share a single power strip. The laser drivers, LB box, function generator (connected to LB box), photodiodes supply, oscilloscope, spectrum analyzer (HP 8560E), and moku all are now powered from the long power strip ('A') located under the optics table. The OCXO preamplifier box supplying ~30 MHz modulations for PDH is powered by the Sorensen DC supplies... which are not powered from the same strip. The Sorensens are on the Aux rack, while the other electronics are on the PSOMA rack. It would be somewhat involved to power the Sorensens from the same supply (need a longer power cable sent across the walkway), but can be done.
  • I tied the optical table to ground through a long (springy) wire with 1 MOhm resistor soldered inline near the halfway point. One end of the wire is bolted to the optical table, and the other is bolted to the enclosure frame. There are two power strips bolted to the enclosure... but the 'enclosure' power strips are plugged in to the wall on the opposite side of the room from the 'under the table' power strip powering the PDH electronics. That's not great. I could move the PDH electronics over to the 'enclosure' power strips, but it would require some cable rerouting.

notes

  • The DB9 cable connecting the North TEC driver to the laser is flaky (moving the cable causes the TEC alarm to trip on the driver). The rubber of the cable is visibly separated from the casing, and the ground shielding is exposed no. I have powered down the North laser, and suggest we don't run it until replacing or repairing the cable
  2833   Fri Dec 10 11:30:44 2021 shrutiUpdateLaserBeat between the lasers

I powered on the 'North laser' using an ITC 510, and make set up the fiber optics for the beat note as depicted in Attachment 1:

  • Powered the Newfocus 1611 FC (fiber-coupled). The rack mount box (Thor Labs RBX32) does not yet have an adapter for the PD power supply so I wired it through an open slot.
  • There are 90-10 beam splitters after the faraday isolators of both the lasers with the 10% coupled output going to a 50-50 beam splitter to form the beat note.
  • The transmission of the 50-50 beam splitter is sent to the 1611-FC
  • I'm not sure how to dump these beams appropriately within the fiber box: the 90% output in the path of the north laser and the anti-symmetric port of the 50-50 BS. Currently, they are just being dumped into metal/ rubber covers.
  • Still cannot find the beat on the spectrum analyzer (either the connections on the 50-50 BS are weird, or I have not properly explored the relative setting space to get a beat within 1 GHz]

Locking updates:

  • The LB box seems to also output some modulating current when the lock is turned off ~ 3mA pk-pk change which is weird
  • I've added a 10dB attenuator at the input of the LB1005 servo box to make sure the PDH signal is within 100 mA (the input limit of the LB1005)
  • I experimented with changing the attenuation at the output at input but realized that without a 10-20 dB attenuation the current ends up being too high in the loop and rails. Currently there is a 10dB attenuator at the output.
  • The output (after attenuation) is now also split with a T to be sent to the Moku for troubleshooting and offset correction
  • I've been trying unsuccessfully to get data from the scope through a LAN connection

I'm collecting useful calibration/gain parameters in Elog 2830 for quick reference.

Attachment 1: beat.pdf
beat.pdf
Attachment 2: fiberbox.pdf
fiberbox.pdf
  2832   Tue Dec 7 11:44:27 2021 ranaNoise HuntingLaserfinding a stable lock
  1. check to see if the 140 kHz noise is on the laser already. Post spectrum of beatnote in ELOG.
  2. Check error signal of cavity with a beam block placed inside the cavity. This allows the reflected beam to still hit the PDH PD to check for electronics pickup, but there is no cavity signal.
  3. Check error signal with beam to REFL PD blocked. also post to ELOG.
  4. Let's discuss about the 50 Ohm termination stuff. Generally, the 50 Ohm or not issue is not critical in these applications where the signal is mostly less than 2 MHz (after demodulation); the wavelength for 2 MHz in cable is ~(c*2/3)/(2 MHz) ~ 100 m. The main purpose of checking impedances is to make sure that the sources can drive the load, whether its 1 Mohm, 50 ohm, or 25 ohm. So it would be fine to have a T going to a 1 Mohm scope input.
  2831   Tue Dec 7 11:11:37 2021 aaronNoise HuntingLaserfinding a stable lock

[aaron, shruti]

Shruti has observed a 144 kHz oscillation on the PDH error signal from our cavity-with-cantilever. The oscillation was railing the PDH signal, making it impossible to maintain lock for more than a few seconds. I came in to troubleshoot.

  1. Observed a flaky lock with PDH signal railing
    • initial settings on the LB box were LFGL = 90 dB, G = 6.76, PI corner = Int
  2. Adjusted gain on the LB box to G=5, then back to 6.76 after observing no change
  3. The PDH signal was being split with a BNC T junction after the lowpass, sending one end to the LB box channel A input (50 Ohm) and the other to the oscilloscope (1 MOhm). I removed the T junction and sent the PDH signal directly to the LB box, and instead sent the LB box's 'Error Monitor' output to the oscilloscope (still 1 MOhm). I observed no significant change
  4. I noticed that the current drive output of the LB box was being split with two T-junctions, each terminated with 50 Ohm, resulting in the LB box's output driving 25 Ohm. I removed the unnecessary T junction, so the output is now sent to a 20 dB attenuator, followed by a T-junction with one end terminated in 50 Ohm and the other sent to the current drive mod in.
    • The cavity can now remain locked for about 10s
    • After reducing the gain to G=5.70 and tweaking the laser temperature, the PDH error signal is not quite railing (oscillating at 142 kHz with 500 mV pkpk)
  5. I realized I should have changed the input impedance for the oscilloscope channel monitoring the PDH error signal in step 3. 'Error Monitor' output wants to drive 50 Ohm.
    • After adding adding a T-junction to 50 Ohms between the error mon output and the 1 MOhm scope, I observed 300 mV pkpk oscillations at 142 kHz and stable locks for up to a minute.
  6. I removed some extraneous BNC T-junctions (were just open ended), and disconnected aux out from the back of the LB box. No change in the 142 kHz oscillation.
  7. I next checked the LB box offsets by
    • turning off the servo and terminating both inputs with 50 Ohm.
    • Tuned the input offset to null the LB box error monitor point (viewed on Moku scope) with the servo off.
    • Also tuned the sweep offset to null the LB box output with the servo off, though this should not matter for the in-loop behavior    
    • No change for the 142 kHz oscillation

That's all for me this morning. I think the oscillation is sufficiently low we could try using the DAQ to feed back to temperature as we were before. It would be useful for diagnostic purposes to maintain a more extended lock, and I'm finding I need to tune temperature anytime I reacquire lock. Maybe we're just always sitting close to the edge of the current control loop.

  2830   Thu Dec 2 13:40:57 2021 shrutiUpdatePSOMAProgress and cavity parameters/ lock parameters

WIP

Videos of flashes and a very noisy lock.

In Attachment 1 and the video showing a lock, the PDH signal essentially goes to its rails. Most of the noise is in oscillations that are roughly at 143 kHz.

 

Cavity and Lock Parameters

This elog (2815) shows the cantilever with its mount and a top view of the cavity.

Optics

  • FSR: ~ 600 MHz
  • Finesse: [Input coupler transmittance: <4% , Cantilever transmittance:]. Check polarization - measure Finesse for both.
  • cavity pole: also depends on polarization
  • DC optical gain (measured as the PDH slope): W/m, W/Hz

 

Electronics

  • Laser resistance->temperature conversion chart
  • Laser temp->frequency to be measured using beat (has not been done consistently in the past since it varies with current)
  • Actuation Coefficient
  • Signal path calibration (first should optimize demod phase)
    • New Focus 1811 40,000 V/A, ~0.85 A/W ?
    • Mixer conv loss: 5 dB
    • 50 Ohm term before LP (6 dB attenuation)
    • compare the W/m calibration to expectation based on:
      • input laser power
      • RF modulation depth
      • cavity Finesse
  • LB lock box transfer function with & w/o Boost engaged.
  • Useful dynamic ranges
    • LB box input stage: +- 10 V. The input stage is a difference amplifier buffer, allowing +- 10 V common mode signals to be subtracted before the filter stage
    • LB box filter stage: +- 330 mV
    • LB box summing stage: +- 10 V
    • LB box output drive: +- 20 mA or +- 10 V
    • ITC 502 current modulation input: +- 10 V
  • Measure transfer functions of laser
    • Measure ITC current driver separately, then ITC+laser
    • Measure phase response using beat and phase lock

 

  2829   Wed Dec 1 13:21:54 2021 shrutiUpdatePSOMATowards locking PSOMA with cantilever

In order to improve the acoustic isolation of the box I used rubber tubing cut as depicted in Attachment 1 and attached to the edge of the plastic box (as seen in Attachment 2).

Attachment 2 also shows one of the four dog clamps I used to secure the box. Finally I placed an oscilloscope on top of the box to provide some weight.

The box has 4 holes for the input, reflected, transmitted-to-PD, and transmitted-to-camera beams.

 

PDH signal

In elog 2731, the PDH signal is almost 1 V pk-pk, but recently I had only been seeing a very noisy signal ~10 mV pk-pk (purple trace in Attachment 4. Transmitted light is yellow, reflected light is blue). I adjusted the lengths and removed the attenuators in the path of the EOM but it did not seem to change.

While going through the path of the cables once more, I realized that signal from the OCXO was being sent to the wrong port on the rack mount (Attachment 5). It was earlier on D3, and I later changed it to A4 which was labelled 'S EOM', but the actual location was D4 which I identified from this diagram. Now I see a reasonably large PDH signal as expected with ~800mV pk-pk (video here).

Attachment 1: IMG_1250.png
IMG_1250.png
Attachment 2: IMG_1252.png
IMG_1252.png
Attachment 3: IMG_1251.png
IMG_1251.png
Attachment 4: IMG_1253.png
IMG_1253.png
Attachment 5: IMG_1254.png
IMG_1254.png
  2828   Thu Nov 18 11:09:24 2021 shrutiUpdatePSOMATowards locking PSOMA with cantilever

Since clamping the cantilever, I began to align the triangular cavity with the cantilever.

Attachment 1 shows the setup of the cavity with the input coupler at M1, curved mirror at M2, and a cantilever with a flat optic at M3. The cantilever does not possess any knobs for fine adjustment and its initial mounting is fixed. M1 and M2 can be adjusted, but to get to TEM00 resonance, and any fine alignment, the two steering mirrors before the cavity were used.

Weird things happen while aligning:

1. While aligning, I began to see resonant peaks corresponding to the cantilever's mechanical resonance showing that the cantilever has a very high Q. I mistakenly thought the cavity was close to being aligned to resonance in this configuration while also being very stable, but later understood that the strong oscillations actually seemed more like accidental misalignments creating an optical lever. The green traces in the videos are the measured transmitted light through the cantilever. When the cavity is really aligned close to the TEM00 mode the transmission looks like the green trace here, and the reflection (from M1) is the blue trace.

2. Another time I thought it was aligned I was looking at the camera at the transmission of the cantilever [M3 in Attachment 1] and could see a single flashing spot but when I moved the camera around I actually ended up seeing two beams instead, simultaneously flashing. Minor changes in the alignment resulted in the beams individually and simultaneously turning into higher order modes. Initially I thought that the situation was as depicted in Attachment 2 and found that it was a consistent geometric solution, but later Koji pointed out that was the geometry that creates the TEM01 mode. In Attachment 2, the initial beam (red, M1->M2) overlaps not with the beam in that path from the second round trip (yellow, M1->M2) but with the one from the third round trip. Was it possible that this misalignment was too severe that instead of being TEM01 two separate, but identical, resonant cavities were created? A measurement of the FSR, which I did not do, would have proved that that was the case.

Since the transfer of fiber components into the box, I have re-aligned the cavity to the TEM00 mode. Note: replacing the fiber at the fiber launch and also changing the polarization at the fiber launch moves the beam around.

PDH locking stuff:

[more information and data coming soon]

1. The PDH signal seems too low, possibly because the LO and RF relative phase must be re-adjusted.

Attachment 1: CavitySetup.pdf
CavitySetup.pdf
Attachment 2: Geometry01.jpeg
Geometry01.jpeg
  2827   Wed Nov 10 10:45:22 2021 aaronUpdateSi fabKOH etching

We've been learning the various processes required to fabricate silicon cantilevers following Zach's recipe.

The last step involves etching cantilevers out of the silicon wafer using KOH. Specifically, Zach's etch recipe for a 500 um wafer coated with 400 nm SiNx on both sides is:

  1. Submerge wafer in 30% KOH solution at 80 C for 6 hours
  2. Remove and rinse the wafer, then scribe and break along the lines between the cantilevers
  3. Return the separated pieces to the KOH bath for an additional 2 hours
  4. Rinse, dry, and briefly HF etch the resulting cantilevers

something wrong with the KOH etch

I tried to follow this recipe last Wednesday, with the following modifications:

  • Our wafer is only 300 um thick
  • KNI now stocks 50% KOH instead of 30% KOH. This is probably good for us, since higher molarity KOH results in more specular etched surfaces. But it does change the etch rate
    • According to Fig 2.28 in Silicon Micromachining (fig from Seidel et al), The etch rate for Si in KOH should decrease from 50 um/hour in 30% KOH to about 30 um/hour in 50% KOH

Because the wafer is 3/5 the thickness but the etch rate is also 3/5 as fast, I anticipated that a 6 hour etch would be appropriate to produce something for scribing and breaking... however, after 4.5 hours, the entire wafer was almost entirely dissolved. All that remained were thin, fragile sheets of Si or SiNx. What's going on? Some possibilities

possible explanations

Expected rate of etching was off

I have since perused the KNI wiki for more resources on etch rates through Si. The most extensive data seem to be from the BYU page on KOH etching. It suggests that 30% KOH at 80 C etches through 100 Si at 80 um/hour (compared to 60 um/hour that Zach was using), while 50% KOH etches through 100 Si at 45 um/hour. This doesn't really explain the results. Even at 45 um/hour, we should have been left with 100 um of Si after 4.5 hours, or 1/3 of the initial material. If we take Zach's 60 um/hour at face value and apply the data from BYU as a 'relative rate of etching,' we would be scaling by a factor pretty close to that suggested in Silicon Micromachining

I wasn't able to find good data on the rate of KOH etching through SiNx depending on temperature or concentration.

We are using doped silicon, but I found reference online to boron doped silicon etching more slowly.

The bath temperature could also have been systematically higher than 80 C, or just not well controlled around 80 C.

To test this, I'm planning to do a shorter etch of a sacrificial piece of Si under the same conditions as before (30% KOH, 80 C). I'll remove any oxide layer with HF beforehand, and check the bath temperature with a thermometer. 

We didn't really have 400 nm of hard SiNx on both sides of the wafer

I used the standard SiNx deposition recipe on both sides of the wafer, but did not check the resulting mask with an ellipsometer. We should in the future do this after most previous steps: PECVD, optical lithography, DRIE etching, etc.

I'm being trained on the ellipsometer this afternoon, and plan to measure the thickness of SiNx on some 4" wafers Zach had left over. 

Wafer was thinner than 300um

Our cleaning process involves a few minutes in an HF bath, and we use DRIE etching during the optical lithography step. Either of these processes could thin the wafer. In particular, during lithography I noticed that my photoresist was a bit thinner than I'd intended. Perhaps the I etched through the exposed photoresist more quickly than anticiapted, allowing the DRIE etch to reach the underlying silicon for longer. 

In the future I'll measure the thickness of the edges of the wafer (where there is no cantilever) with a micrometer before etching.

 

Update

This afternoon, I tried to measure the etch rate of the KOH bath. I did the following:

  • Auto-tuned the PID parameters of the KOH bath's temperature controller. The settings ended up being unchanged from before
  • Heated the bath to a nominal temperature of 80 C. I put a thermometer in the bath and found the true temperature to be 85 C. After finishing for the day, I set the calibration offset of the temperature controller such that the thermometer reads the true temperature. 
  • Tested the rate of etching with a wafer I'd recently RCA cleaned
    • The wafer is from the same batch as the previous wafer: 3" diameter, boron doped with resistivity 1-10 Ohm-cm, and double side polished
    • Before starting, I used a micrometer to measure the thickness of the wafer at 0.30 mm. 
    • Prepared a solution of 2.5% HF, and etched the wafer for 2 minutes then rinsed with DI water
      • After the HF etch, I measured the thickness of the wafer at 0.29 mm. The purpose of the HF etch was to remove any SiO2 from the surface of the wafer before bathing in KOH.
    • Place the wafer in KOH bath (85 C true temperature) for 35 minutes. The wafer was completely dissolved in this time.

I had intended to remove the wafer and measure its thickness again, so unfortunately can only place a lower bound on the etch rate. Nonetheless, the implied rate of etching is >16.5 um/minute (etched through at least 290 um in 35 minutes, from both sides of the wafer). This is more than an order of magnitude faster than expected, even allowing for the increased bath temperature. Clearly I am missing something -- is the KOH actually being diluted due to some additional DI water in the pumping system? Is the boron doping really increasing the etch rate by that much? Did the wafer just fall off of its holder and get lost in the murky KOH (I did fish around for several minutes, no sign of a wafer in the bath)?

  2826   Tue Nov 9 19:01:43 2021 aaronUpdateLaserfiber components to fiber box

I mounted the fiber components (north and south Rio lasers, Faraday isolators, 90-10 beamsplitters, 50-50 beamsplitter, fiber EOM, 1611FC-AC) inside our breadboard-in-a-box from Thorlabs (attachment 1). 

Along the way

  • Inspected and cleaned all fiber tips that I connected. Fibers that are not connected have not necessarily been cleaned. I took some photos using my phone and a 15x macro lens (attachment 2 is of a clean fiber), but the camera kept focusing on some schmutz on the microscope's eyepiece, so they're not very illuminating. 
  • Turned off the S laser that Shruti was using, disconnected its cables, and reconnected the cables through the DB9 feedthrough on the box
  • Removed from cryo cavs table the beamsplitter, photodiode, and fiber connectors that we were borrowing for the three corner hat measurement. These are now inside the box.
  • labelled all fiber ends
  • Used the 10m patch cable that was running from cryo cavs table to PSOMA table to carry the beam from the fiber box to the beam launch on the PSOMA table. Note that one end of the patch cable had dust or damage on its face over the fiber core that I couldn't manage to clean off. Could try again, or get another (shorter) patch cable). 

Though I set up the beamsplitters in the box approximately where we will eventually want them, the beam currently does not pass through any beamsplitters. I left the optical layout for the south laser identical to how it was on the table, with the addition of a patch cable between the EOM and the launch: laser -> faraday -> EOM -> patch cable -> launch. The north laser is not connected to its drivers. 

Attachment 1: IMG_2187.jpg
IMG_2187.jpg
Attachment 2: IMG_2184.jpg
IMG_2184.jpg
  2825   Tue Oct 26 10:01:54 2021 shrutiUpdatePSOMATowards locking PSOMA with cantilever

26 Oct 21, Tue

Yesterday while the power was out I turned off laser drivers and other powered electronics that I could think of. When the power was back I rebooted the computers (checked that I was able to ssh into cymac) and the electronics that I had turned off. I noted that the cavity was back on resonance as indicated by the forest of peaks that we were seeing.

Today when I returned to the lab the forest was missing and Aaron noticed that the cantilever seemed to be tilted in the clamp. Attachment 1 shows the exact angle at which it was found. A small part at the end is chipped off but it otherwise appears usable to me.

Attachment 1: The cantilever at the angle at which it had tilted to with the chipped bits, after just removing the top piece of the clamp.

 

Before re-clamping, we (Aaron and I) decided to use a variable torque socket wrench to test for optimal clamping torques on another sacrificial cantilever. I decided to abandon this experiment after Rana pointed out at the group meeting that this would not give any really reliable number for the required clamping torque since breakage torques may have a wide range of random values for different cantilevers.

 

28 Oct 21, Thu

Attachment 2: Mobile phone image (without external lens) of the cantilever's clamping surface. It shows some pits and scratches, one larger than 100 microns.

For reference, the dimensions of the clamping area are 1cm x 1cm.

 

Attachment 3: The clamping surface of the smaller (top, while clamping) piece of the clamp, that contains the indentation where the cantilever sits.

Attachment 1: TiltedCantilever.pdf
TiltedCantilever.pdf
Attachment 2: cantileversurface.pdf
cantileversurface.pdf
Attachment 3: clampsurface.pdf
clampsurface.pdf
  2824   Wed Oct 20 11:31:06 2021 shrutiUpdatePSOMATowards locking PSOMA with cantilever

[aaron, shruti]

Aaron removed the plastic wrap and began powering and reconnecting all electronics.

I drilled three holes into the plastic box that covers the cavity, reconnected the south laser diode and TEC to the ITC 502 combi controller and set the laser temperature to 9.888 kOhms and current to 126 mA (until the beam was visible).

 

21-Oct-21

Continued with alignment of the cavity until I observed a forest of peaks again while sinuoisoidally oscillating the temperature, although they can be observed even without temperature cycling now that we ahve a cantilever.

 

Next steps:

  • Mode-match to TEM00 better to be able to observe something on the IR camera
  • Drill holes for both the transmission PD (behind cantilever) and transmission monitor (behind curved mirror)
  • Set up locking electronics once again
  • Weigh and clamp down the plastic box
  • Procure new mount for the cube PBS to mount it in path so s-polarization transmits
  • Add PBS and HWP to the input path
  • Set up beat with another free-running laser to get a sense of the laser noise when measuring the noise spectra
  2823   Tue Oct 19 10:45:08 2021 aaronLab InfrastructurePSOMAPSOMA table enclosure assembly

[aaron, shruti, raj]

We added the acrylic framing that supports the HEPA FFU, as well as two of the 3 panels covering the top of the enclosure. Raj and Shruti also unwrapped and added handles to the orange acrylic doors.

However, the last acrylic panel for the roof of the enclosure is too small. Though the part is cut as specified at 23" on the short side, the counterbore holes on the long side of the panel do not span the distance between the two rails where the panel mounts. The counterbore holes are 0.75" from the edge of the panel, which means the two lines of holes are 21.5" apart. However, the distance between the center of the rails is specified at 22.5". The distance between the rails is appropriate for the size of the HEPA FFU, and the 0.75" gap along the edge of the panel is standard across the other roof panels. Therefore, I suspect the overal width of the panel was just specified 1" too short.

I'll alert F&L and have them send another panel ASAP

 

  2822   Thu Oct 14 09:02:57 2021 aaronLab InfrastructurePSOMAPSOMA table enclosure assembly

[aaron, shruti, chris, raj, radhika]

On Tuesday, Shruti and I did fit checks of all connectors for the enclosure. We received the remaining parts according to the parts list from F&L, so should have everything we need. We requested clarification on where to use a few connectors, though haven't yet received a reply (our contact at F&L was unavailable).

  1. Add the feet to the vertical bars (E)
  2. Connect cross struts to the horizontal bars (U and V to G and T)
  3. Slide horizontal bars (G and T) onto the vertical bars (E), starting with the short side (G)
  4. Add the plate brackets to the framing bars
  5. Connect the framing on the top of the enclosure (A) to the horizontal bars along the top of the enclosure (B)
  6. Connect the remaining framing (F and A) along the top of the enclosure, first with the end connectors then with brackets
  7. Add the plastic panels to the top of the enclosure
  8. Add acrylic doors

With some adjustments, we completed steps 1-5. The frame of the enclosure is around the table. Tomorrow we'll try to complete the rest of the build, which includes adding the roof and panel doors to the enclosure.

  2821   Wed Sep 29 10:27:50 2021 ChrisComputingstuff happensbackups

Backups were restarted for the cryo lab computers gaston, spirou, and cominaux. A 4TB USB drive was connected to cominaux, mounted under /backup, and rsnapshot was configured to run on a nightly basis. It does not back up the full disk, but only those directories where user-generated files are kept (/home, /etc, /usr/local, /opt, /ligo). rsnapshot's configuration files are: /etc/rsnapshot.conf and /etc/cron.d/rsnapshot.

For the cymac, configuration and minute trend files are backed up by rsnapshot to a 1TB disk, mounted as /backup on cymac1.

  2820   Wed Sep 29 09:57:50 2021 aaronLab InfrastructurePSOMAPSOMA table enclosure assembly

[shruti, aaron, raj, chris, ian]

  1. Prepping
    1. Powered off lasers and other electronics around the table, disconnected power cords, and separated the PSOMA rack several inches from the table
    2. Wrapped the table in plastic in preparation for assembly
  2. Assembling the frame
    1. Brought in the 80-20 bars, and set the long 2x1 bars along the sides of the table ('x' is 95" long, 'y' is 50", 'z' is 90.25")
    2. Identified the enclosure's feet, which are PN #65-2191 (bolt with a flat hex head). We're pretty sure from the enclosure diagram that these feet mate with the rectangular block PN #2130. Unfortunately, the bolt is M10 while the threaded block is 3/8-16... happily, we found some other rubberized feet with a 3/8-16 bolt in the EE shop, and were able to mount those on the vertical bars.
    3. Next up, we noticed that none of the 45 degree cut cross bars have arrived. As far as I can tell, we have all other parts, but I have

I've asked F&L for a parts list so we can do a proper inventory before starting assembly. Also requested a quote on purple panels, update on shipment for the cross bars, and imperial threaded feet.

  2819   Tue Sep 28 15:47:39 2021 aaronLab InfrastructurePSOMAPSOMA table enclosure

Shruti and I transferred the materials for our new enclosure from receiving to the subbasement hallway. I noticed that the acrylic panels we received are amber, not purple as we specified... unfortunately it looks like the error was already present in the quote, which listed acrylic #2422 (in my emails with their rep, we'd both confirmed acrylic #2424). Because it was in the quote, I'm not sure there's anything we can do, but I'll ask. 

Ian, Raj, Shruti, Chris, Aaron, and possibly other grads will meet at 10 tomorrow for assembly.

  2818   Wed Sep 22 11:12:12 2021 PacoMiscEquipment LoanReturn lens

We returned CFC-2X-C into the same cabinet it was borrowed from. Thanks CRYO!

Quote:

I've borrowed 2 collimating lenses (for fiber input) for use with the free-space AOM in the DOPO lab.

 

  2817   Tue Sep 21 10:53:01 2021 shrutiDailyProgressPSOMAin-air cantilever cavity

Today I began to see a forest of peaks in the transmission (transmission through the cantilever optic) while aligning meaning that the cavity is around resonance. I then adjusted the Watec camera to the transmission of the curved mirror and began to see some flashes. I took a video showing this which can be found here.

I tried activating the feedback loop with the settings we had on earlier but while it did seem to increase the peak powers of the peaks it did not seem to lock. The steering mirror to the refl photodiode (Newfocus 1811) and the corresponding lens needs to be re-adjusted.

 

 

  2816   Mon Sep 20 16:31:37 2021 shrutiNoise HuntingPSOMAtfs and noise before changing out optic->cantilever

Similar to elog 2767, before changing out the mirror from a rigid mounted one to the cantilever, I measured the loop transfer functions and noise spectra since we had not done so since we moved the table.

The only change between the current and previous version was that I tried to make the low frequency phase offset zero.

Attachment 1 shows different estimated open loop transfer functions corresponding to ratios of specified closed loop transfer functions (as mentioned in the labels). The UGF was lower than what was previously acheived.

Attachment 2 shows the fit of the slope of the open loop transfer function as done previously in elog 2776.

Unfortunately, I did not notice the issue with binning when I measured the noise spectra which resulted in a discontinuous spectra corresponding to the different regions I measured separately on the Moku. The data for this and everything else mentioned in this post is in Attachment 3.

 

Attachment 1: OLTF.pdf
OLTF.pdf
Attachment 2: OLTF_fit.pdf
OLTF_fit.pdf
Attachment 3: Data20210915.zip
  2815   Mon Sep 20 14:08:18 2021 shrutiUpdatePSOMAin-air cantilever cavity

In short: Between Friday and today, I de-bonded the mirror from the broken cantilever, picked two cantilevers and bonded both the mirrors to them with the AR surface being the one on the bond side, clamped and mounted on a post one of these cantilevers, and placed it in the cavity replacing one of the flat mirrors (M3).

1. De-bonding from broken cantilever

For this, I used methanol and soaked it for around 30 min similar to the procedure here. Since this bond was weaker, it took lesser time.

2. Bonding both mirrors to cantilevers

I picked one cantilever from the previous collection and another from the dish shown in Attachment 1. This second cantilever did not seem to have the back surface passivated with SiNx and looked like it was just oxidized silicon.

I used the GE cryo-varnish at four points around the mirror placed on the etched edge of the cantilever for both and let it dry for a few hours (Attachment 2). The bonding was performed such that the AR surface was on the side of the cantilever.

3. Clamping and aligning

Using the alignment jig on the table, I picked the cantilever without the SiNx passivation on both sides (identifiable as with one non-glossy side) to mount to the clamp. I tightened it not too strongly and then used a 3/4" post of suitable height to get the center of the mirror to a height 4" from the table.

I placed it roughly where I thought it should be on the table replacing the flat mirror M3 from the previous setup. I got the beams from two round-trips to overlap visually and added a PDA10CS to look at the transmission from the cantilever.

The aim now is to do as we did earlier - sinusoidally change the temperature by driving the TEMP TUNE input with a function generator and slowly tweak the alignment of the steering mirrors to find the cavity resonance.

 

I also measured the transmission of the cantilever as roughly 0.1% using the power meter.

Attachment 1: source.pdf
source.pdf
Attachment 2: twobondedcantilevers.pdf
twobondedcantilevers.pdf
Attachment 3: clampingalign.pdf
clampingalign.pdf
Attachment 4: clamped.pdf
clamped.pdf
Attachment 5: cantileverincav.pdf
cantileverincav.pdf
Attachment 6: cantileverincav1.pdf
cantileverincav1.pdf
  2814   Thu Sep 16 19:15:31 2021 aaronDailyProgressstuff happensanother mirror bonded to another cantilever

I broke the cantilever while fastening the fork clamp. Afterwards, I used methanol to remove the varnish and separate the mirror from the cantilever shard. I then used a cotton-tipped swab soaked in methanol to clean the varnish from the sides of the mirror. I drag wiped the HR and AR surfaces of the mirror with methanol followed by isopropyl alcohol. Finally, I bonding the mirror to a different cantilever (this one with somewhat more pitting than the previous) -- again bonding with cryo varnish at four points on the sides of the mirror, but on recommendation from Chris this time with the mirror AR surface touching the cantilever.

  2813   Wed Sep 15 16:45:54 2021 aaronDailyProgressOpticscantilever clamped

This morning, I clamped the cantilever-with-mirror that we bonded yesterday.

I first used a broken cantilever to practice the clamping. I clamped the test piece first in the nominal position, then with slight alignment errors in the available degrees of freedom to see what those errors would look like. If the cantilever is misaligned such that it does not rest in the clamping groove, the clamping block will show a larger than usual gap on at least one side (attachments 1, 2). If the cantilever is not entirely 'in' the clamp, the thicker part of the cantilever will be visible above the clamp (attachment 3). Attachment 4 and 5 show the test cantilever in good alignment.

Next, I clamped the cantilever-with-mirror by 

  1. Insert the alignment pins into the clamping block
  2. Place the bottom clamping block flush with the aligning piece
  3. Place the cantilever on top of the clamping block and aligning piece, such that its long edge is flush with the aligning edge and the thick part of the clamped Si is just at the end of the clamping block
  4. Set the top clamping block into place, holding the bottom block in place and lining up with the alignment pins. Make sure the gap between the two clamping blocks is even, and that the thinned part of the cantilever ends just at the edge of the clamping blocks.  
  5. Add the clamping screws and tighten enough to secure the cantilever. Attachment 6 is the final result. 

The Q may improve with more clamping force, but I'd like to do a more controlled test of how much torque can safely be applied to the clamping screws. We don't want to break any usable cantilevers. 

Attachment 1: gap_side.jpg
gap_side.jpg
Attachment 2: gap_bot.jpg
gap_bot.jpg
Attachment 3: protrude_above.jpg
protrude_above.jpg
Attachment 4: goodClamp_above.jpg
goodClamp_above.jpg
Attachment 5: goodClamp_bot.jpg
goodClamp_bot.jpg
Attachment 6: clamped.jpg
clamped.jpg
  2812   Wed Sep 15 16:42:31 2021 RadhikaMiscEquipment LoanBorrowing lenses for DOPO lab

I've borrowed 2 collimating lenses (for fiber input) for use with the free-space AOM in the DOPO lab.

Attachment 1: IMG_2532.jpeg
IMG_2532.jpeg
  2811   Tue Sep 14 16:43:11 2021 shrutiUpdateOpticsDebonding and bonding mirrors to silicon cantilevers

[Shruti, Koji, Aaron]

Today Koji guided us to (1) remove the mirror I had contacted to the cantilever that later broke, and (2) clean and contact the second flat (1550 nm wavelength coated) mirror to another cantilever using cryo varnish.

(1) De-bonding the mirror

After my failed attempts at using methanol and acetone along the edges followed by soaking the contacted-to-silicon optic in isopropanol, today we were finally able to de-bond the mirror by soaking in methanol. We used two washers to raise the optic from the aluminum surface and folded foil pieces to keep everything in place. The setup shown in Attachment 1 was covered with another aluminum foil dish and let sit. It took at least an hour of soaking to completely de-bond.

Koji then drag wiped the optic with pure isopropanol to clean any remaining residue and cryo-varnish hairs.

 

(2) Bonding a different mirror to a cantilever

First we tested drag wiping the broken silicon cantilever with methanol and isopropanol. Then we selected a reasonably looking cantilever that was not chipped at the edges, though it had an uneven surface, and drag-wiped both surfaces with isopropanol.

We drag-wiped the HR surface of the optic a few times and then placed it on the portion of the cantilever that had the square etched-out region. While Koji held the mirror in place I applied cryo-varnish to four points around the mirror. This is now set to dry.

 

 

Update: Photos are available on the ligo.wbridge google drive.  I uploaded everything, but could pare down to save drive space. 

Attachment 1: varnishdissolve.pdf
varnishdissolve.pdf
  2810   Wed Sep 8 09:11:43 2021 RadhikaMiscEquipment LoanBorrowed 1550nm AOM

I borrowed this free-space AOM (1550 nm) for use in the DOPO lab.

Attachment 1: IMG_2518.jpeg
IMG_2518.jpeg
Attachment 2: IMG_2517.jpeg
IMG_2517.jpeg
  2809   Tue Sep 7 10:01:39 2021 shrutiUpdatePSOMAattempt at bonding, another broken cantilever...

Attachments 1 and 2:

The cantilever clamp with a cantilever on it to be used on an optical post. This will replace the non-transmitting flat mirror in the PSOMA cavity. There are two screw holes to mount the optical post, but it would only be possible to use one, therefore this will be initially mounted asymmetrically.

Attachment 3:

Cantilever with the two flat mirrors from Zach's setup. Without having it properly mounted, I could only very crudely measure its reflectivity as R>81% using a power meter.

The one on the left is shown with (what I believe, by comparing to the curved optic in our present setup, is) its HR surface facing up and the one on the right with its AR coated surface shown.

 

Attachment 4:

I attempted to bond the mirror to the cantilever with cryo-varnish after testing that it bonded two pieces of silicon.  The HR surface which would be facing the interior of the cavity was the side I decided to put the varnish on because the pictures from Zach's thesis depicted the cantilevers that way. I coated the edge of the square-shaped hole on the cantilever head with the varnish and placed the mirror on it. I placed a lens wipe on top of the mirror and then another flat optic over it. While applying pressure on it with a tweezer the cantilever broke (possibly because the aluminium foil below it was very crinkly which I didn't think too much about beforehand).

While the mirror seems to have bonded after a couple of minutes of holding it down, I probably have to remove the varnish (with IPA/acetone I think ?) and re-bond to a different cantilever.

Attachment 1: CAA96DEB-236A-4BD4-B196-8985374562AB.jpeg
CAA96DEB-236A-4BD4-B196-8985374562AB.jpeg
Attachment 2: 03813DE5-AC9D-4958-A203-60B0C8751BCC.jpeg
03813DE5-AC9D-4958-A203-60B0C8751BCC.jpeg
Attachment 3: CA57BFB3-9A8C-4AAF-ADA0-F11F507D8464.jpeg
CA57BFB3-9A8C-4AAF-ADA0-F11F507D8464.jpeg
Attachment 4: B458F0DC-8BD8-422B-9F21-88CF18163C0B.jpeg
B458F0DC-8BD8-422B-9F21-88CF18163C0B.jpeg
  2808   Wed Sep 1 12:20:18 2021 ranaDailyProgressLaserquieting the Rio W and Teraxion beat notes, but not much help

What about showing the E/W beat noise for several different operating points on the W laser? Temperature and current.

  2807   Tue Aug 31 15:14:28 2021 aaronDailyProgressLaserquieting the Rio W and Teraxion beat notes, but not much help

Between 0.5 and 50 Hz, there are a couple of regions where the Rio W laser noise dominates the three corner hat measurement. And, below 0.3 Hz, the Teraxion laser noise dominates the measurement. Today I'm going to try to quiet these two lasers a bit to make a slightly improved three corner hat estimate.

I'll be looking at the output of the delay line frequency discriminator (DFD) on the moku spectrum analyzer, so I've swapped in the 1.9 MHz lowpass filter for the one Rana was using to check the noise out at MHz.

I saved the following traces on an SR785, but with a nonfunctioning GPIB so ended up storing them on a floppy drive.

  • Baseline Rio W x Teraxion beat note spectrum, as measured by phasemeter to Moku / SR785
  • Beat note spectrum after securing the fibers and Faraday isolators to the table, and moving the Rio W laser to a different operating point. 
  • "" after adding a layer of foam atop the fibers
  • "" after turning off the HEPA blowers. Noticed several noise lines around 100-300 Hz substantially reduced after turning off the HEPA.

I saw minor improvements to the Rio W x Teraxion beat note spectrum, and took 20 minutes of data for each beat note. Dropbox upload failed several times, so I sent it to my laptop via ipad file storage and airdrop. After swapping fiber connectors and covering with foam, the system took ~15-30 minutes to equilibrate each time (though perhaps longer would have been prudent, since I still saw low frequency drift up to 4 MHz during the 20 min measurement time). No improvement to the estimate on Teraxion laser, and if anything the West laser was even more noisy relative to Rio E and Teraxion, across a wider frequency band (almost the entire band from 30 mHz - 90 Hz). The foam and turning off the HEPA FFU did reduce the noise below 1 Hz, especially for the Rio E x Teraxion beat note. Figure 1 uses maximal averaging for every 2-fold frequency increase, on 20 minutes of data taken today (attachment 2 reproduces the relevant figure from last week's data, without the Marconi reference for better viewing). 

Attachment 3 is the updated Teraxion noise estimate. 

Update: attachment 4 is a comparison of the frequency noise (uncalibrated) for three different configurations: after taping down the fibers, after adding foam, and after turning off the HEPA blowers. I'm not sure why turning off the HEPA blowers increased the noise, maybe should have let the system settle longer? Despite the overall higher noise floor with HEPA blowers off, several peaks between 10 and 100 Hz were reduced. 

Attachment 1: beats_regions_base2.pdf
beats_regions_base2.pdf
Attachment 2: beats_regions_base2.pdf
beats_regions_base2.pdf
Attachment 3: Teraxion_noise_base2.jpg
Teraxion_noise_base2.jpg
Attachment 4: DLFD_foam.pdf
DLFD_foam.pdf
  2806   Tue Aug 31 13:23:05 2021 aaronDailyProgressSuspensionclamped cantilever

[aaron, shruti]

We found a heavily pitted cantilever from Zach's early fab runs, and mounted it in his clamp using the alignment jig and pins. We tried optically contacting one of the mirrors to the cantilever, but some combination of surface oxidation and roughness (or just inexperience) prevented us from making a bond. The surface of the cantilever had obvious defects, so we weren't very hopeful. We're seeing what Koji suggests for glue.

 

  2805   Mon Aug 30 14:06:00 2021 aaronNotesSi fabCantilever fabrication process

We're starting to make new cantilevers this week. Here is our process, largely drawn from Zach's thesis, based on the process from the Chao group (D1200849) and the standard techniques of hard mask etching.

Wafer stock

Zach started with a 100 mm (4") undoped <100> Si wafer with 500 um thickness. He reports achieving similar mechanical Qs of the final cantilevers when starting with either SSP or DSP polished wafers, though we may want to investigate this further. To avoid spoiling too many large wafers, I'd like to start by processing our 2" x 280 um wafers that are leftover from the cryo Q experiment, and will inventory our larger wafers to determine what we should order. We also have some 3" wafers we can use to fabricate full-length (7 cm) cantilevers in smaller batches than had we used 4" wafer. 

Process

  1. RCA clean, following the steps in elog 2343 but with the HF etch after the RCA2 bath. 
  2. Deposit a 400 nm thick hard mask of SiNx with PECVD
  3. Pattern the nitride mask with photolithogrphy
    1. Apply 1.5 um of AZ 5214E photoresist evenly on one side of the wafer, using a spinner run at 4000 rpm for 1 minute
    2. Bake the wafer with photoresist on a hot plate at 110 C for 50 seconds
    3. Load the wafer and a photolithography mask into the Suss Microtech MA6/BA6 contact aligner, which exposes the photoresist to UV for ~ 10 seconds
    4. Develop the photoresist by bathing in MF CD-26 for 1 minute, then rinse and dry the wafer
  4. Dry etch the nitride with a pseudo Bosch etch, which uses a plasma of sulfur hexafluoride and octafluorocyclobutane. The process produces a vertical etch profile by passivating the exposed surfaces during the etch. 
    1. Etch for 5 minutes to remove the 400 nm SiNx layer in the exposed region
    2. Then, use a remover solution (chemical not mentioned) to remove the remaining photoresist.
  5. Wet etch through the remaining Si in the exposed region
    1. Submerge the wafer in a 30% KOH solution at 80 C. At this temperature, the etch proceeds at 1 um / min, so it takes ~ 8 hours to etch through a 500 um wafer.
    2. 6 hours into the etch (for a 500 um wafer), remove the wafer from the KOH bath and break the remaining wafer into rectangular pieces along the etched lines
    3. Place the rectangular pieces into a fixture and return them to the KOH bath for the remainder of the etch.
    4. Rinse, dry, and finally etch away the remaining SiNx hard mask by submerging briefly in HF
  6. Finally, passivate the surface by depositing a thin (10s nm) layer of SiNx with PECVD

Zach's procedure calls for thinning the central region of the cantilever, which softens the suspension and improves isolation. I expect we'll want to evaluate our cantilevers before thinning, since the procedure to thin the central region is a bit tricky. When we do thin the cantilevers, we will not terminate the process at step 6 but instead continue with the following:

  1. Deposit another SiNx hard mask
    1. deposit 200-300 nm SiNx on the 'bottom' side of the cantilever with PECVD
    2. Also deposit 200-300 nm SiNx on the 'top' side of the cantilever, but physically obstruct the central region of the cantilever with a sacrificial piece of Si. Covering part of the cantilever ensures SiNx is mostly deposited only on the exposed regions.
  2. Another round of wet etching
    1. Etch in KOH for 4 hours, which leaves about 250 um thick region in the center of the cantilever (versus 500 um on the ends). 
  3. HF etch the remaining SiNx, DI rinse, and dry the cantilevers
    • It's unclear to me whether to passivate with SiNx after the final HF etch, or if the final etch is simply terminated before the bare Si surface can be exposed. Will ask Zach.
    • Following the final HF etch, package the cantilevers in a wafer carrier and seal in a plastic bag purged with dry N2 gas.  

Materials

Materials we need to acquire are in bold

  • RCA clean
    • Chemicals provided by KNI
    • Wafer from our stock
    • PVDF reinforced wafer tweezers and baskets for use with the HF bath
  • PECVD
    • Possibly a better wafer holder, but the standard cassettes will do for now
  • photolithography
    • Possibly need to acquire our own developer, does KNI carry MF CD-26?
    • We found what we think is the appropriate transparency mask in Zach's old materials
  • Etching
    • KNI stock materials only

Equipment

Our training status on the equipment is green for 'full user,' orange for 'supervised user,' or red for 'need initial training.

  • Oxford Instruments System 100 PECVD unit
  • Oxford Instruments System 100 ICP 380 etcher
  • Suss Microtech MA6/BA6 contact aligner
  2804   Fri Aug 27 15:27:38 2021 ranaDailyProgressLaserdelay line frequency discriminator

rather than the perfect solution to PSD estimates, how about using the code you already have and just change the binning a little more often than once per decade? i.e. stich together as you already did, but get more averaging in the noisy spots. Should be a very easy modification to the code.

  Draft   Thu Aug 26 19:40:06 2021 ranaNoise HuntingLaserhigh frequency frequency noise measurement w/ DFD and new low pass

I modified Aaron's DFD box: I replaced the SLP-1.9 (1.9 MHz low pass filter) with a 44 MHz low pass that we conveniently had in the lab.

The purpose of this is to measure the high frequency noise of some of our lasers to figure out if there's any difference between the RIO Planex and Teraxion NLL.

I did the hookup and everything looks good so far. So far I have measured the RIO E/W beat and the RIO E vs Terax beat.

Need to do some more data processing to get plots.

Some notes along the way:

  • For monitoring the output of the RF amp, we want to use the CPL IN port of the bi-directional coupler that goes between the amp and the DFD. The CPL IN measures a coupled version of the input, whereas the CPL OUT measures the reflection that comes back from the DFD box. At 175.5 MHz, with the 44 MHz low pass on the mixer IF port, the reflected signal ~ -40 dBm when the CPL IN was at ~ -10 dBm.
  • I have made a shared folder in ligo.wbridge dropbox called PSOMA_data. Now that Moku is logged in to dropbox, and dropbox installed on the gaston workstation, the files are automatically synced to the workstations. We can also share the folder with our personal dropbox accounts so that there's seamless data availability.
  • Also attaching the times series plot used to calibrate the discriminant. I used the Marconi knob to scan the frequency and find the min & max of the DFD box at

Attaching beat note noise spectra showing the noise of the DFD (with no pre-amp, so probably Moku input noise), the RIO Planex (E & W lasers from the Cryocav table), and the Teraxion NLL. Looks like they're similar in HF noise.

Attachment 1: TeraRIOEast_beat_Screenshot.png
TeraRIOEast_beat_Screenshot.png
Attachment 2: Planex-Beatnote-ASD.pdf
Planex-Beatnote-ASD.pdf
Attachment 3: DFD-Discriminant-Cal.pdf
DFD-Discriminant-Cal.pdf
  2802   Thu Aug 26 11:58:50 2021 aaronDailyProgressLaserdelay line frequency discriminator

Yesterday (August 25), I measured another 10 min at 488 Hz of Moku phasemeter data for the three pairs of beat notes. All three lasers had been on overnight, so there was no longer low frequency drift of the Teraxion laser. Rather than amplifying then picking off the beat note, I sent the RF output of the 1611 directly to the Moku's phasemeter input.

what kind of spectral density estimate to use?

Today, I've been figuring out how to get more averages out of our data. One approach (the one used above) is a modification of Welch's method:

  1. Start Welch's method with as large a window as possible, given the desired number of averages. For example, if there are 2**16 data points and we want at least 4 averages, then use Welch's method with 50% overlap and 2**15 points per segment. 
  2. From the first application of Welch's method, save the first N frequency bins (including the DC value). Our spectral estimate runs from 0 to (N-1)*f_0 in steps of size f_0.
  3. Next, decrease the number of samples per segment by a factor of N, and repeat Welch's method. In our example above, we are now able to take 4*N averages. Assuming the rounding is handled correctly, the frequency resolution of the second Welch's periodogram is f_1=N * f_0. 
  4. From the second estimate, save the N-1 samples from f_1 to (N-1)*f_1.
  5. Continue applying Welch's method to successively smaller window sizes and saving the 'low frequency' data from each. The iteration terminates when the window size is unnacceptably small (for example, nperseg <= 2), at which point you can save the remaining spectrum up to the Nyquist frequency.

The above procedure sacrifices some frequency resolution at the higher frequencies in exchange for additional averaging. The tradeoff with resolution is necessary, because the window size determines not only the smallest resolvable frequency, but also the spacing of frequencies in the spectrum. For the spectra from the previous elog in this thread where N=10, the total measurement time is 10 minutes, and the sampling rate 488 Hz, there is evidently more noise higher in the frequency decade (7-9e^n) than lower (1-3e^n). More consistent averaging can be achieved by setting N=2, but at the expense of most of the high frequency resolution (only 33 frequency bins survive the procedure). 

One workaround is to modify the procedure so the frequency binning is mostly set at the beginning, by the 'highest resolution' available. Then, perform Welch's method with as small a window as possible while still resolving the frequency bins. Care must be taken at high frequency: eventually, the 'FSR' of our Welch's method cannot resolve an f_0 difference in frequency into an integer change in the number of samples per segment. The spectrum can either be cut off at that frequency, or the procedure can continue while accepting nonstandard bin widths at high frequency.

Other workarounds are perhaps less desirable. One could accept nonuniform frequency binning, and simply compute Welch's method for every available choice of nperseg. This would maximize the number of averages in each bin, but especially at low frequency, there will be substantial correlation between adjacent frequency bins. Another workaround is to save the entire spectrum wherever evaluated, then combine the data later. One must again worry about correlations between the measurements: at high frequency, we would be combining coarse data with many averages with fine data with fewer averages. 

Another approach entirely is to do something smarter than Welch's method. In our meeting today, Chris suggested I look into multitapering. Spectral estimates can reduce bias due to leakage by introducing a tapered window, at the cost of increased measurement variance. Welch's method heals the variance relative to standard tapering by overlapping the windowed segments, at the cost of some frequency resolution. Multitapering instead minimizes loss of information by increasing the number of degrees of freedom of the estimates. The Here are a few resources on the topic:

While checking out Percival and Walden, I stumbled across parametric methods for spectral estimation -- those where an early spectral estimate is used to refine the procedure and spectral estimate iteratively. Perhaps up the alley of some recent discussions at our group meeting.

error propagation

Simply applying Gaussian error propagation is not quite right, because the PSD is exponential distributed (the ASD is Rayleigh distributed, see Evan's note T1500300). Each ASD is Rayleigh distributed with

  • mode \sigma
  • mean \left(\frac{\pi}{2}\right)^{1/2}\sigma \equiv \mu
  • rms \sqrt{2}\sigma = \frac{2}{\sqrt{\pi}}\mu 
  • median \sqrt{\ln 4}\sigma
  • variance of \frac{4-\pi}{2}\sigma^2 \equiv \delta(X_{ij})^2

For a large number N of averages, the central limit theorem lets us estimate the mean of each PSD, x_{ij}^2, with normal distributed uncertainty and variance \delta(X_{ij}^2)^2/N. Our three corner hat estimates of the ASD Y are based on the scaled, root mean-squared sum of three such PSD estimates, so for each frequency bin we can estimate the variance of the laser's ASD by

Y = \frac{X_{ab}^2 + X_{ac}^2 - X_{bc}^2}{2} \\ \delta(X_{ij}^2)=2x_{ij}\delta(X_{ij}) \\ \delta(X_{ab}^2 + X_{ac}^2 - X_{bc}^2) = \sqrt{\delta(X_{ab}^2)^2+\delta(X_{ac}^2)^2 + \delta(X_{bc}^2)^2} \\ \implies \delta(Y)^2 = x_{ab}^2\delta(X_{ab})^2+x_{bc}^2\delta(X_{bc})^2+x_{ac}^2\delta(X_{ac})^2\\

I'll use the final equation above, along with the number of averages, to estimate the uncertainty in each frequency bin of the final frequency noise ASD of the individual lasers. In particular, the filled region is \sqrt{Y \pm \delta(Y)}

Attachments

[OK, I'm having a lot of trouble uploading pdfs to the elog this week, even with rasterizing. I've dropped these figures along with one set for the case of 'Welch's with no averaging' onto gaston under /home/controls/cryo_lab/Figures/3CH ]

  1. Time series of Rio E x Rio W beat frequency
  2. "" Rio E x Teraxion ""
  3. "" Rio W x Teraxion ""
  4. "" Marconi ""
  5. Welch's estimate of the frequency noise on the beat notes, with increased averages every decade
  6. Three corner hat estimate of the frequency noise on the Teraxion laser, from the ASD in (5)
  7. Welch's estimate of the frequency noise ont he beat notes, with increased averages every factor of 4 in frequency
  8. Three corner hat estimate of the frequency noise on the Teraxion laser, from the ASD in (7)
  9. Welch's estimate of the frequency noise on the beat notes, with averages increasing every factor of 2 in frequency
  10. Three corner hat estimate of the frequency noise on the Teraxion laser, from the ASD in (9)

I think these results warrant a more careful measurement, especially in the decade around 1 Hz. Also, the error bars are obviously way underestimated. 

Attachment 1: EW_timeseries.pdf
EW_timeseries.pdf
Attachment 2: EX_timeseries.pdf
EX_timeseries.pdf
Attachment 3: WX_timeseries.pdf
WX_timeseries.pdf
Attachment 4: Ma_timeseries.pdf
Ma_timeseries.pdf
Attachment 5: beats_regions_base10.pdf
beats_regions_base10.pdf
Attachment 6: Teraxion_noise_base10.jpg
Teraxion_noise_base10.jpg
Attachment 7: beats_regions_base4.pdf
beats_regions_base4.pdf
Attachment 8: beats_regions_base2.pdf
beats_regions_base2.pdf
Attachment 9: Teraxion_noise_base2.jpg
Teraxion_noise_base2.jpg
  2801   Tue Aug 24 16:44:39 2021 aaronDailyProgressLaserdelay line frequency discriminator

[aaron, rana]

  • Drive RF amplifier with -6 dBm from Marconi (amp is +16 dB), and observe RF level at the 1% RF couple out just before the delay line box
  • Sweep the Marconi carrier frequency from 75 to 230 MHz at 1 MHz / 50 ms. Observed a somewhat assymetrical sine wave, indicating there's some switching or other issue giving us amplitude-to-phase coupling at the mixer (see Rana technical docs above)
  • Removed the Marconi drive and power off RF amp, then open up the delay line box to add a 4 dB attenuator on the RF path.
  • The output of the delay line box is now a more or less symmetrical sine wave across our sweep
    • Turned off the sweep, and tuned the carrier frequency to record a 382 mV pkpk output of the discriminator between about 75 MHz (higher voltage) and 145 MHz (lower voltage). THe null is around 108 MHz
  • Disconnected the Marconi from the system, and instead plugged in the Rio laser E x Rio laser W beat note from the 1611. Tuned the laser current to null the discriminator output, and recorded the beat frequency with the Moku phasemeter. The lasers have been on for a while and are no longer drifting, so apparently it's staying within the phasemeter bandwidth.
    • After measuring Rio E x Rio W, swapped in the Teraxion laser and measured Rio E x Teraxion. The Teraxion laser exhibited some low frequency drift (several MHz / min), which seemed to improve over time
    • Then, measured Rio W x Teraxion
    • Lastly, I sent a sine wave from the Marconi (amplitude chosen such that the signal going into the moku remained -35 dBm) into the phasemeter to measure its noise floor.

Data logged:

File name measurement notes
EW_note_20210824_164439.li Rio E x Rio W laser, near 112 MHz, with Moku phasemeter. 488 Hz sampling  
EW_note_20210824_170434.li Rio E x Rio W laser, near 112 MHz, with Moku phasemeter. 15.6 kHz sampling  
EX_note_20210824_175820.li Rio E x Teraxion laser monotonic, low frequency drift, several MHz / min
WX_note_20210824_184820.li Rio W x Teraxion laser  
Marconi_20210824_194102.li Pure sine at 110 MHz  

We uploaded the data from moku to dropbox, and pulled it to spirou via the web interface. Should give the workstations dropbox.

Later, I realized we could probably have done better by sending the output of the 1611 directly to the moku. Instead, we were amplifying +16 dB then using an RF coupler to pick off 1% for the moku.

Figures:

  1. ASD for the three beat notes, using 10 minutes at 488 Hz sampling for each. Each curve is the ASD from Welch's method with bias-adjusted median averaging, 50% overlap and a Hanning window. The shaded regions are the 15.8-84.1% percentiles. Welch's method is applied 'decade-by-decade', so within each frequency decade the window size is adjusted to maximize the number of averages (as in labutils/moku/modifiedPSD.py). The ASD of the Marconi sine wave is also shown for a rough noise floor reference.
  2. The ASD for the Teraxion laser using a three corner hat with the PSD from attachment 1.
  3. The remaining figures are the time series plots, for reference.
Attachment 1: beats_regions.pdf
beats_regions.pdf
Attachment 2: Teraxion_noise.pdf
Teraxion_noise.pdf
Attachment 3: EW_timeseries.pdf
EW_timeseries.pdf
  2800   Tue Aug 24 16:25:17 2021 ranaElectronicsGeneralapp note on mixers vs phase detectors

https://www.minicircuits.com/app/AN41-001.pdf

https://www.markimicrowave.com/blog/all-about-mixers-as-phase-detectors/

https://www.rfcafe.com/references/articles/wj-tech-notes/Mixers_phase_detectors.pdf

  2799   Mon Aug 23 14:35:36 2021 shrutiDailyProgressPSOMAre-aligning

[shruti, aaron]

After plugging back in all the cables, I was struggling to get the cavity locked again or seeing the transmitted beam on the monitor.

Today, following Aaron's suggestion, I removed the connection from the PDH mixer IF port to the oscilloscope and also the power splitter I had added earlier (now sending the IF output directly to the LB input only with the 10 dB attenuator). The cavity didn't lock right away but changing the gain to 6 and re-adjusting the input and output offset got the cavity locked again.

When we tried to get it on the monitor, we saw that the mode it locked to was a higher order mode misaligned in the vertical direction. We first aligned the beam vertically until we started seeing lower order modes and finally the TEM00 mode.

I adjusted the lenses a little to tweak the mode-matching after aligning the beam into the cavity for maximum transmission when locked. Calculating the mode-matching using the reflected beam when locked and unlocked shows that it is only 60% though.

 

  2798   Mon Aug 23 13:38:57 2021 aaronDailyProgressLaserdelay line frequency discriminator

cymac ADC noise

The noise I was seeing last week on the ADC did not show up when driving the same channels directly with a function generator, only when buffering the function through the SR560. Wrapping the BNC several times around a ferrite toroid between the SR560 and ADC reduces the noise to close to the level of the ADC noise floor (there is a < 5 count pkpk sine wave cross-coupled into the adjacent channels for a ~2000 count pkpk sine wave on the channel of interest, but the signal carrying channel itself looks clean).

The SR560 that was overloading on battery last week is now also overloading on line power. I've swapped it with one of our functioning SR560.

calibration

  • A 113.2 MHz sine from the Marconi nulls the output of the frequency discriminator, as measured by the G=1, DC-coupled SR560 sent to an oscilloscope.
    • At 113.2 MHz, a 314.2 Hz FM with 800 kHz deviation appears as a 1.24 mV sine wave on the G=10, AC-coupled SR560 (measured by an oscilloscope).
    • The DC-coupled SR560 has a 1 kHz lowpass filter, the AC-coupled has a 3 kHz lowpass filter (both 6 dB/oct)
  • I measured the noise level of the Marconi over several 10s of minutes (while trying to figure out how to use the calibration features of diaggui). The result is in attachment 1, with the y-axis still in units of 'counts / rtHz'.

delay line

Afterwards, I looked at the spectrum for the Rio E x Rio W laser beat note. Looked OK... there's some kind of filtering happening in my delay line though. If I watch the spectrum of the RF coupler's pickoff (1% between the amplifier and delay line box), the implied peak power entering the delay line box is ~ 5 dBm near any of the nulls (so ~113 MHz or 176 MHz), but over 15 dBm between the nulls. I hadn't noticed this behavior before (with the busted mixer and RF coupler before the amplifier).

Attachment 1: Screenshot_from_2021-08-23_16-42-48.png
Screenshot_from_2021-08-23_16-42-48.png
  2797   Fri Aug 20 11:03:44 2021 aaronDailyProgressLaserdelay line frequency discriminator

I'm driving the input of the delay line box with 10 dBm from Marconi's RF output.

  • First, tuned the Marconi carrier frequency to 112.8 MHz to null the output of the delay line mixer (after DC coupled, G=1, dynamic range mode, SR560 buffer with 1 kHz, 6 dBm/oct lowpass).
  • Next, frequency modulating the RF carrier at 314.14 Hz with 800 kHz deviation from the Marconi. The output of the delay line box is an 16 mVpp sine wave, see attachment 1
  • I checked out the noise spectrum of the Marconi buffered through both the DC coupled SR560 above, and separately an AC-coupled, G=100, low noise mode SR560 with 10 kHz, 6 dBm/oct lowpass
    • On the Moku, the spectra look reasonable. The G=1 and G=100 peaks at 314 Hz differ by 40 dB, as expected. Attachment 2.
    • Feeding the signal to our ADC has some issues. Not sure what's going wrong (ground loops? hot swapped AA-ADC cable?), but any signal I plug in to the ADC results in the periodic noise in attachment 3 on not only the signal-carrying channel but also the adjacent channels. The discontinuities are larger than the G=1 buffered signal, though not as large at the G=100 buffered signal.
      • I'm using the PSOMA channels, so the names are innaccurate. In attachment 3, the DC-coupled signal is in SLD_PDH_CTL, while the AC-coupled signal is in SLD_PDH_SIG
      • The noise is at about 120 Hz, which suggests some power line issue. Measuring only the AC-coupled SR560 in battery powered mode, with no other instruments connected after the buffer, improves but does not eliminate the noise (attachment 4). Also, on battery powered mode, the SR560 constantly overloads; my exact sequence of steps to observe this overload behavior was (1) power off the SR560, (2) unplug the line voltage, (3) power on the SR560.
Attachment 1: 30A43669-5DFF-400A-BEB4-7EF71BBAE1A6.jpeg
30A43669-5DFF-400A-BEB4-7EF71BBAE1A6.jpeg
Attachment 2: A3BF5DA5-32E5-42A5-A1D8-9270C3D88518.png
A3BF5DA5-32E5-42A5-A1D8-9270C3D88518.png
Attachment 3: Screenshot_from_2021-08-20_16-11-26.png
Screenshot_from_2021-08-20_16-11-26.png
Attachment 4: Screenshot_from_2021-08-20_16-20-48.png
Screenshot_from_2021-08-20_16-20-48.png
  2796   Thu Aug 19 16:32:47 2021 ranaDailyProgressLaserdelay line frequency discriminator

Ack!! stay out of the "Fix the SR560s" gamecool Do not repair, do not send back.

Let's get a trustworthy measurement of the frequency noise ASAP.angel

Quote:
 

I've borrowed one SR560 from CTN to get us by for now, but am wondering if I should order parts to make repairs, or simply send the units to SRS?

 

  2795   Thu Aug 19 14:21:35 2021 aaronDailyProgressLaserdelay line frequency discriminator

I replaced the mixer from my delay line box with a functioning ZFM-2-S+. I also replaced the lossy pink SMA cable (122 cm) with a slightly longer (150 cm) cable that I made from a spool of LCOM coax. I also replaced the shorter cables with solder soaked ones. The final build is in attachment 1.

I drove the delay line frequency discriminator with a swept sine from the Marconi, and observed a symmetric response in the IF output. Sweeping from 5 MHz to 80 MHz is in attachment 2, showing that the response is symmetric about zero from -107 to 108 mA.

I'll return later this afternoon for a more careful calibration and to take some data. 

Update, calibration

I'm calibrating the delay line discriminator by running a swept sine from the Marconi, and this time reading back on an acromag channel.

Attachment 3 is the first run (y axis in Volts), where the Marconi is ramping from 1 MHz to 240 Mhz, stepping by 10 kHz every 100 ms. There are some surprises. The output is linear until about 70 MHz, possibly because the mixer is only good down to 5 MHz. The upper half fringe is not symmetric to the lower half fringe, even comparing the curve only above 10 MHz.

I'm repeating the measurement with different sweep parameters: 1 MHz to 250 MHz, stepping 1 kHz every 50 ms. The steps will be somewhat faster than the Acromag's sampling rate, but those units dither so it should smooth out. I'll post some plots with the time axis converted to frequency when the data come in.

Attachment 1: 007FC2D7-F29B-4636-9CE7-04CBC3B789F7.jpeg
007FC2D7-F29B-4636-9CE7-04CBC3B789F7.jpeg
Attachment 2: 1F1E68F0-199F-4E1C-B909-883A521E06CF.jpeg
1F1E68F0-199F-4E1C-B909-883A521E06CF.jpeg
Attachment 3: Screenshot_from_2021-08-19_17-53-58.png
Screenshot_from_2021-08-19_17-53-58.png
  2794   Wed Aug 18 09:55:13 2021 aaronDailyProgressLaserdelay line frequency discriminator

I'm taking this data today, here are the issues I've hit

  • Period noise around 100 Hz visible in the output of SR560.
    • Replaced with a different SR560, this noise is gone
  • 3.7 MHz harmonics around the beat note. I was seeing these before, changing laser temperature and current don't affect these spurious peaks. I'll look for them if I measure the laser current noise again.
    • My guess was I'm saturating the 1611 PD. The beat note power is almost the full 1 mW linear range of the device. However, reducing the beat note power to -10 dBm does not eliminate the harmonics, especially the first.
  • First SR785 I used was getting stuck interfacing via GPIB, it's labelled
  • I'm calibrating the delay line by T-ing off the signal after the lowpass and sending it to a DC-coupled SR560 with G=1, then an oscilloscope. When I take spectra, I'll record the AC and DC coupled SR560 ouputs on AC- and DC-coupled input channels of the SR785 respectively.
Laser pair beat note power E laser TEC setpoint E laser LD current W laser TEC setpoint W laser LD current lower null frequency lower null DC voltage upper null frequency upper null DC voltage beat frequency at start beat frequency at end Hz/V cailbration Timestamp notes
cryo cavs E x W -5 dBm 11.738 kOhm 56.56 mA 8.146 kOhm 54.22 mA 120.2 MHz -215 mV 199.9 MHz 32 mV 158.0 MHz     Wed Aug 18 12:08:57 2021 calibrated during second measurement
cryo cavs -2 dBm 11.738 kOhm

55.88 mA

8.145 kOhm 53.24 mA 12.02 MHz -215 mV 199.9 MHz 32 mV 159.7 MHz 162.6   Wed Aug 18 15:37:37 2021 still at end of linear range on 1611

 

SR560s in need of repair

We have 7x SR560 in need of repair in the cryo lab. Most of these (at least 5) are constantly overloading, which indicates the front end transistor and op amp need replacing. From SRS, these parts are $50 each, but I found I think the same parts on digikey (the FET transistor is LSK389B; the amp is OP37A) for about $10 each. There are at least 2 that need a new battery.

I've borrowed one SR560 from CTN to get us by for now, but am wondering if I should order parts to make repairs, or simply send the units to SRS?

mixer busted

Rana noticed that my delay line is asymmetrical -- sweeping the frequency through a full cycle of the interferometer should reverse the sign of the output, but instead the output was biased negative.

Indeed, when I drive the mixer RF with 0.5 Vpp at 25 MHz from a function generator while supplying the LO with 7 dBm at 150 MHz from the Marconi (see attachment 1), my ZFM-3-S+ has an asymmetric waveform despite at most a few mV difference in the LO and RF DC levels (see the video in attachment 2). The same test with a different mixer (ZFM-1-S+) gives the balanced waveform in attachment 3.

I'll replace the mixer, and make a few modifications to the cables of my delay line box.

Attachment 1: 0A96102F-5465-456B-93BF-E2B9D4A7E96C.jpeg
0A96102F-5465-456B-93BF-E2B9D4A7E96C.jpeg
Attachment 2: trim.2553D656-98A0-442D-9185-4C5252B2B727.MOV
Attachment 3: trim.787F04FC-8DED-4501-B67D-FF98EC5D5D56.MOV
  2793   Tue Aug 17 14:04:27 2021 aaronDailyProgressLaserdelay line frequency discriminator

I'm repeating this measurement, but moving the photodiode and electronics to the cryo cavs table, so we don't send anything over a long, floppy fiber across the room. See attachment 1 for diagram.

I set up the measurement, but didn't take any data. What's the right way to find the IP address for these GPIB controllers? I ended up scanning with nmap, but suppose I should mess with prologix' netfinder tools to assign our controllers static IP addresses and just label them.

Attachment 1: 3A1E242C-F4F6-4A87-B2E8-E74839BE542C.jpeg
3A1E242C-F4F6-4A87-B2E8-E74839BE542C.jpeg
  2792   Mon Aug 16 18:48:37 2021 ranaDailyProgressVacuumventing cantilevers cryostat

could the pumpdown plot be made so that the units are visible? Maybe use dataviewer or python?

 

  2791   Thu Aug 12 15:03:32 2021 shrutiDailyProgressVacuumventing cantilevers cryostat

[aaron, shruti]

We repeated steps 1-4 in elog 2789 and, with two people, managed to get the o-ring to stay in place while lowering.

  • Before pumping down the cryostat, we decided to pump down everything until the valve to the cryostat (by closing the valve to the cryostat), as a check. In around 6 min the pressure dropped to 100 utorr and kept decreasing to tens of utorr slowly. After 60 min we stopped this experiment.
  • We then opened the valve to the cryostat and began pumping down, the pump stopped automatically after around 10 min. This is the first kink in curve in Attachment 1. At this lower pressure, we decided to restart the pump anyway. Even though the turbo came up to 90000 rpm pretty quickly, the lowest pressure achieved after an hour was not lower than 400 utorr.
  • Thinking that the low rate may be because the long tube that connects the cryostat to the pump was of a small diameter, we decided to change the setup to one with the larger diameter. Attachment 2 has the final setup.
  • Then starting again at atmosphere, we began pumping down again. It does seem to go faster, although the pump did stop once and had to be restarted. I think we plan to leave it pumping overnight.
Attachment 1: Screenshot_from_2021-08-12_17-51-06.png
Screenshot_from_2021-08-12_17-51-06.png
Attachment 2: 885A9694-996E-499E-BDCC-FB550338561F.jpeg
885A9694-996E-499E-BDCC-FB550338561F.jpeg
  2790   Wed Aug 11 16:09:29 2021 aaronLab InfrastructureLab Workpreparing for lab rearrangement

During the vacuum testing, I'm continuing some lab maintenance, mostly cleaning up and labeling cables on the electronics racks.

  2789   Wed Aug 11 13:21:23 2021 aaronDailyProgressVacuumventing cantilevers cryostat

The cantilevers cryostat vacuum line [edit: Aaron (May 2022) suspects I was pumping on the vacuum line only, not the chamber, as indicated in the previous log] has only reached 7 utorr (the reading on the gauge matches epics) after pumping overnight. I'm going to try manually actuating the gas ballast valve in case our roughing pump is manual-only.

  1. Turned off the pumping station, and wait for turbo to spin down before venting up to air.
  2. Manually move the roughing pump's gas ballast valve from the 'open' to the 'closed' position, as in figure 2 of the MVP 015-2 manual (roughing pump). I spent some time digging through the HiPace 80 (turbo pump) manual and figuring out how the venting valve on the side of the turbo works.
  3. Turn on the pumping station and observe the pumpdown. The result is in attachment 1.

After closing the gas ballast valve, the pressure drops below 7 utorr in under 30 minutes and is approaching 1 utorr. That's not the best I've seen from this pump, but should be good enough to continue diagnosing the cryostat.

To that end, I'm installing a new 2-270 Viton o-ring on the cantilever cryostat (needs name).

  1. Clean the new o-ring with isopropyl alcohol and a lint-free wipe
  2. Grease the new o-ring with Krytox, and place on a fresh sheet of HV aluminum foil
  3. Use the crane to open the cryosta at the midsection. The screws have already been removed.
  4. Install the o-ring and close the cryostat, tightening the bolts in a star pattern
    • I couldn't get the o-ring to stay in the inverted groove. I found Zach mention he taped some L-brackets to hold it in place, but couldn't find L-brackets nor could get our Kapton tape to stick to the cryostat with the weight of the o-ring. Will try it again with two people.
Attachment 1: Screenshot_from_2021-08-12_14-57-42.png
Screenshot_from_2021-08-12_14-57-42.png
  2788   Tue Aug 10 17:39:40 2021 aaronLab InfrastructureLab Workpreparing for lab rearrangement

Friday, 6 August

Facilities came to move out the cryo Q (central, rigid legs) optics table, and left us the legs. They also took the two desks from the lab. Lastly, they moved the PSOMA table to the center of the room and rotated it 90 degrees so its long axis runs EW. Afterwards, Shruti and I moved the PSOMA rack to the E end of the PSOMA table, directly across from the other two racks .

Tuesday, 10 August

  • set up new desks
  • installed gaston and spirou workstations on the desks
  • Began rerouting cables among the three electronics racks.
    • We would like to move the AI and AA chassis over to the cymac rack, to reduce the number of cables running between racks and get them away from the DC supplies. Has the added benefit of making them somewhat more accessible in the new configuration. Is there any obvious reason we would not want the Acromag, AI, or AA chassis on the backside of the cymac rack?
    • Continued labeling cables as we go
  • power on Sorensen, Acromag chassis for pumpdown measurement.
  • Removed plastic wrap from the PSOMA table
  • We should figure out what to do with the lHe cryostat. If it leaves the lab, it would be convenient to move the silicon / fiber / optoelectronics cabinet slightly south, to let us space out the desks and place the speakers along the N wall (and easier access for that cabinet).

Attachment 1 is a screenshot of the temperature and particle count trends over the last 4 weeks. I need to figure out how to add axis labels on ndscope... the units for both temprature channels are F, relative humidity is %, and all particle count channels are log(counts). There is a factor of 10-1000 excess particle counts around the time of our lab rearrangement, mostly affecting the larger particle sizes. Humidity also experienced a mild increase. The AD590 temperature channel drops off because I turned off its power supply to accommodate our rearrangement, and haven't yet turned it on again. The daily and weekly particle count fluctuations are interesting, presumably dictated by lab access.

Attachment 1: Screenshot_from_2021-08-10_19-01-13.png
Screenshot_from_2021-08-10_19-01-13.png
  2787   Tue Aug 10 17:23:39 2021 aaronDailyProgressVacuumventing cantilevers cryostat

[aaron, shruti]

Today we are testing the pumping station while pumping on just some blanked off Ts and the vacuum gauge. Last time, we pumped on the vacuum hose leading to the closed Key valve at the cryostat, and observed the pressure level off near mtorr, before eventually reaching only several utorr. The turbo should really have no trouble getting to utorr pumping on just hose sections, so we'll try to observe some better pumping action today.

  • Powered on acromag chassis and set up pressure gauge on just the turbo pump as in attachment 1
    • Noticed that at atmosphere, the gauge reads 744 torr, but the epics channel reports 722 torr.
    • Pumped down on the blanked off system, and again observed the pressure level off at 20-30 utorr after 10 min. If we let it continue, perhaps would reach a similar pressure of several utorr. The discontinuity around 10 mtorr is due to the gauge changing modes.
    • Vented the system by valving off the turbo then using the up-to-air valve go to atmosphere. You'll note the vent looks a little faster than intended; the valve had been too zealously tightened.
  • Following the instructions in the HiCube Eco manual for troubleshooting a pump that does not reach desired final pressures or takes a long time to get there, we suspect either condensation has built up in the gas ballast line or the gas ballast line has been left open. Aaron went through these troubleshooting steps back in 2019 a couple times. We need to open then close the gas ballast line.
    • Following the manual and my procedures from 2019, we first closed the red vent valve on the side of the turbo and confirmed that the black vent valve is also closed
    • Next, we went turn Vent Mode on. We found vent mode was already on, which could explain the mediocre vacuum pressures.
    • Just in case, we pumped down again on the same system in attachment 1, this time with the red valve on the turbo closed, and held vacuum for 20 minutes.
    • Turned off the pump, and wait for the turbo to spin down. After the pump reaches 0 rpm, go up to air (this time, we didn't valve off the turbo and opened the up-to-air valve more slowly).
    • Change the valve mode setting to 'auto'
    • Pump down once more to see if it made any difference.
    • We recorded the pressure during the above operations, see attachment 2. A zoomed in view of the last two pumpdowns is in attachment 3. The final pumpdown was moderately faster, and we've left the pump on to see what final pressure it achieves overnight. Note that we did not re-open the red valve at the side of the turbo; we suspect we probably should have, but aren't sure exactly why. We can try another pumpdown with it open tomorrow.
  • In my previous elog (2289), I noted that our roughing pump appears to have a manually actuated valve on the gas ballast line, and ended up opening and closing that valve. We didn't quite get to try that today, but have noted it for the future.
Attachment 1: 10CC78ED-EC51-4F10-8CBB-9ED236746501.jpeg
10CC78ED-EC51-4F10-8CBB-9ED236746501.jpeg
Attachment 2: Screenshot_from_2021-08-10_19-13-12.png
Screenshot_from_2021-08-10_19-13-12.png
Attachment 3: Screenshot_from_2021-08-10_19-14-19.png
Screenshot_from_2021-08-10_19-14-19.png
ELOG V3.1.3-